A stunt man drives a car at a speed of 20 m/s off a 26-m-high cliff. The road leading to the cliff is inclined upward at an angle of 20∘

Part A:

How far from the base of the cliff does the car land?

Express your answer to two significant figures and include the appropriate units.

Part B:

What is the car's impact speed?

Express your answer to two significant figures and include the appropriate units.
Answer #1

Initial Velocity =20m/s at 20 degrees

so uucos20- 20cos20- 18.79m/s

u,-usin20 = 20-si n 20-6.84m/s

The car will rise a little then falls

Consider the vertical motion of car

Rise

Use Formula v^{2}-u^{2}=2as

v_{y}^{2}-u_{y}^{2}=-2g*y

(Om/s)2 - (6.84m/s)2--2 9.81m/s2 y 6.84m/s) 2 * (at max height velocity becomes zero)

02-6.842ー一2 * 9.81 * y

6.842/19.62 = y

y 2.385m

-------------------

Time taken to rise

v=u+at

v_{y}=u_{y}-gt_{r}

0m/s=6.84m/s-9.81m/s^{2}*t_{r}

0=6.84-9.81t_{r}

{color{Red} t_{r}=0.7s}

----------------------

Falling

Falling Height =y +cliff height =2.385m +26m =28.385m

Time for falling

Use formula s=ut+1/2at^{2}

H=u'_{y}t_{f}+1/2gt_{f}^{2}

Initial velocity is Zero

H=1/2gt_{f}^{2}

28.385m=0.5*9.81m/s^{2}*t_{f}^{2}

28.385=0.5*9.81*t_{f}^{2}

{color{Red} t_{f}=2.41s}

-----------

Total time,t=t_{r}+t_{f}=0.7s+2.41s=3.11s

{color{Red} t=3.11s}

----------

Consider the horizontal motion of car

There is no acceleration in horizontal direction

Speed =Distance /time

Distance=Speed*time

x=u_{x}*t

x=18.79m/s*3.11s

Part A: ANSWER: {color{Red} x=58.44m}

=======================

Part B

u_{x} remains same

so u_{x}=v_{x}=18.79m/s

Vertical motion

Use Formula v^{2}-u^{2}=2as

v_{y}^{2}-u'_{y}^{2}=2gH

v_{y}^{2}=2gH

v_{y}=sqrt{2gH}

v_{y}=sqrt{2*9.81m/s^{2}*28.385m}

v_{y}=23.6m/s

===========

v=sqrt{v_{x}^{2}+v_{y}^{2}}

v=sqrt{(18.79m/s)^{2}+(23.6m/s)^{2}}

ANSWER: {color{Red} v=30.17m/s}

=========================